OpenStudy (anonymous):

For the table below, determine the function rule, and find each of the missing y-values. Function Rule: ? x y 1 4 3 6 4 7 5 ? 7 10 10 ?

OpenStudy (campbell_st):

well you can see that as x increases from 1 to 3 2 places y increases from 4 to 6 2 places so the equation looks like y = x + b you just need to add( or subtract) a number using x = 3 and y = 6 y = 3 + b whats the value for b does it work for x = 1 and y = 4 4 = 1 + b then you have your equation and can dind the missing values.